# Single-Shot Readout and Weak Measurement of a Tin-Vacancy Qubit in Diamond

### v1

@@ C.3 Extended data
C.3.1 Readout fidelity vs. power
Figure 10: Readout fidelity vs. power. Data is fit to Eq. 28.
Studying the power dependence of readout is important for the optimization of fidelity. Notice in Fig. 4b that while the number of signal photons $\bar{n}_{b}-\bar{n}_{d}$ plateaus near saturation, the total number of collected photons continues to increase with power. We attribute this rise to scatter of the excitation laser into the collection path (see Appendix A.1 for details).
This power dependence makes clear that readout fidelity can be optimized by operating near saturation. Below saturation, increasing power simply produces more signal. Above saturation, however, driving with more power does not increase signal and can only cause rising noise. To test this intuition, we plot readout fidelity vs. power in Fig. 10, measured from the same dataset used to produce Fig. 4b. Data are fit to the model in Eq. 28. This model assumes that emitted photons follow Poissonian statistics; see Appendix A.1 for a derivation.
Note that this fit uses $p_{\mathrm{sat}}=313\,\mathrm{nW}$ and $\eta=0.992\times 10^{-3}$ obtained from fits to Fig. 4b, and fits for $f_{0}=0.89\pm 0.032$ and $\chi=977\pm 72$ as free parameters. Fixing $\chi=2244$ as a fixed parameter, as expected from the data in Fig. 2, does not produce a fit which closely matches the data in Fig. 4b. This discrepancy could be attributed to non-Poissonian statistics of light emitted from the ${}^{-}$ (an assumption of Eq. 28), and/or, from detrimental effects due to ${}^{-}$ instability (e.g. cycles where the CRC is passed, but a blinking event occurs before or during the readout step). As such, the models in Eq. 6 and Eq. 28 should be taken as simplifications that do not capture every possible effect on readout of the system, and the parameters reported from these fits should be taken as estimates based on these simplified assumptions, only.
@@
Figure 11: Timing diagrams of the experiments in: (a) Fig. 3, (b) Fig. 4b, (c) Fig. 5. Illustrations are not to scale. Steps are generally separated by $2\,\mathrm{\mu s}$ pauses (not shown). Weak measurement pulses in (b) and (c) resonantly drive the A1 transition, as in the spin polarization and projective measurement steps.
C.3.2 Charge resonance checks
@@
The readout characterization experiment in Fig. 3 includes charge resonance checks (CRCs) applied at both the beginning and end of each characterization cycle. A CRC consists of driving both the A1 and B2 transitions while simultaneously collecting emission. Many photons are collected during the CRC if the excitation is indeed resonant (e.g. neither the laser nor transition frequencies have drifted), and, the system has not left the qubit subspace (e.g. it has not “blinked off”).
@@
The distribution of counts collected during a CRC are shown in Fig. 12a, using data from the “wfm-A” experiment in Fig. 3. Data from wfm-B are nearly identical. The CRC result is recorded as a “pass” if $\geq N_{c}$ counts are measured and a “fail” if $<N_{c}$ counts are measured, where $N_{c}$ is the CRC discrimination threshold. Choosing $N_{c}=30$ for both CRCs for the data in Fig. 3, we find that $7.4\%$ of cycles of wfm-A pass both CRC’s, and $7.7\%$ of cycles for wfm-B pass both CRC’s, out of $4.27\times 10^{5}$ cycles total for each waveform.
@@
The probability that consecutive cycles pass either the first or second check is plotted in Fig. 12b, using a pass threshold of $N_{c}=8$ (the inflection threshold between the dark and bright distributions, Fig. 12a). This probability decays exponentially, consistent with fail events occurring at random. However, probability falls faster for the second check, which may result from drive-induced blinking [35].
@@
Figure 12: (a) Distribution of counts collected during the charge resonance checks (CRCs) in Fig. 3. Each CRC consists of simultaneously driving the A1/B2 transitions. Separate distributions with few or many counts per check are separated by an inflection point around 8 counts. Choosing a pass threshold of $N_{c}=30$ leads to $\approx 15\%$ cycles passing the first check, and of these, $\approx 50\%$ pass the second check. (b) Probability of consecutive passes before a fail is recorded, using $N_{c}=8$. Data are fit to exponential decay. Faster decay is observed for the second check, likely due to drive-induced charging in the intervening readout steps. In both panels only data from wfm-A is plotted; data for wfm-B is nearly identical.
@@
Figure 13: Analysis of the time tagged dataset shown in Fig. 3, as a function of charge resonance check (CRC) threshold photon number $N_{c}$. If both CRCs per cycle each produce a photon number $\geq N_{c}$, the readout steps in that cycle are kept for analysis. (a) Readout fidelity $F_{r}$, and conditional fidelity $F_{c}$ (first readout step and wfm-A, respectively; results from the second readout step and wfm-B look nearly identical). (b) Mean photon number in the bright and dark distributions (first readout step; results from the second step look nearly identical). (c) Probability of passing each CRC, both CRCs, and the probability of the second CRC given a pass of the first. (Data is from wfm-A; results from wfm-B look nearly identical.)
@@ C.3.3 Coherence time
The qubit’s coherence time is measured using a CPMG-2 pulse sequence [64], Fig. 14, operating at $|\vec{B}|=125\,\mathrm{mT}$ and $\zeta=147\degree$. We measure a spin coherence time of $T_{2}^{\mathrm{CPMG-2}}=270\pm 30\,\mathrm{\mu s}$ using a CPMG-2 dynamical decoupling sequence [64]. This measurement is comparable to results in Ref. [26], which operate at $\chi\approx 80$. For spin-1/2 group-IV qubits, at some magnetic field orientations dynamical decoupling has limited effect because the qubit’s $g$-factor is too similar to the predominantly spin-1/2 bath [26]. The measurement in Fig. 14 shows that here we avoid this detrimental effect.
@@ C.3.4 Quantum jumps
We take advantage of single-shot readout to observe real time quantum jumps of the electron spin [42]. The protocol, illustrated in Fig. 15a, consists of a green repump, then spin polarization, followed by readout on the same state that was polarized in the presence of a weak microwave drive. During this readout step, time-tagged photon detection events are binned into $10\,\mathrm{\mu s}$ windows and plotted as a time trace in Fig. 15b. In the absence of the weak spin drive we expect $\approx 0$ photons due to high-fidelity initialization of the spin state and thus a complete dark time-trace of the collected fluorescence. However, the weak spin drive (microwave power $\approx 1\leavevmode\nobreak\ \,\mathrm{dBm}$ at the sample) occasionally kicks the spin-state from $\ket{\uparrow}$ to $\ket{\downarrow}$- for example when initialized using the A1 transition. This results in photons being emitted which can be seen as discrete spikes in the time-trace as our readout determines the state in $\ket{\downarrow}$. Using a photon threshold of $N_{r}\geq 2$ (black dashed line), we distinguish between the “bright” and “dark” state and denote the electron being in the $\ket{\uparrow}$($\ket{\downarrow}$) state. Optical re-initialization also happens naturally under the continuous pump, so the jumping between the spin states (bright and dark) is a mixture of bidirectional mixing (microwave) and one-way reinitialization (bright to dark). To ensure capturing the quantum jumps is not convoluted with charge or spectral hopping, the entire sequence is bracketed with CRCs before and after each measurement shot (using a photon threshold of $N_{c}=30$ for both CRCs). In this experiment, the spin flip process from $T_{1}$ (order Hz at the 1.7 Kelvin sample temperature [26, 27]) is much slower than the readout process. For clarity, Fig. 15c shows the same time trace where the electronic state is assigned as “0” or “1”, based on the discrimination threshold ($N_{r}$).
@@ C.4 List of data
Table 3: Summary of data and operating conditions. Repump power is typically order $100\,\mathrm{\mu W}$ and resonant excitation power is typically order $1-10\,\mathrm{\mu W}$, both specified going into the cryostat. Note that in all data the qubit is driven with the first sideband of the laser created by an electro-optical modulator, not the carrier signal.
Data
Figure
Magnet: amplitude $|\vec{B}|$
Magnet: angle $\zeta$
Repump duration
Init. duration
Readout duration
Cycle time
CRC?
Transitions vs. $\zeta$
Fig. 2c
180 mT
swept
$1\,\mathrm{\mu s}$
$1\,\mathrm{\mu s}$
$22\,\mathrm{\mu s}$
no
Spin polarization
Fig. 2d
180 mT
$53\degree$ and $147\degree$
$100\,\mathrm{\mu s}$
$100\,\mathrm{\mu s}$
$202\,\mathrm{\mu s}$
no
Cyclicity vs. $\zeta$
Fig. 2e
180 mT
swept
$100\,\mathrm{\mu s}$
$100\,\mathrm{\mu s}$
$202\,\mathrm{\mu s}$
no
Rabi
Fig. 2f
125 mT
$147\degree$
$500\,\mathrm{\mu s}$
$100\,\mathrm{\mu s}$
$100\,\mathrm{\mu s}$
$719\,\mathrm{\mu s}$
no
Rabi vs. $\zeta$
Fig. 2g
132-183 mT
$147\degree$
$20-500\,\mathrm{\mu s}$
$5-120\,\mathrm{\mu s}$
$1-60\,\mathrm{\mu s}$
$44-679\,\mathrm{\mu s}$
no
Single-shot readout
Fig. 3
125 mT
$147\degree$
$1000\,\mathrm{\mu s}$
$100\,\mathrm{\mu s}$
$50\,\mathrm{\mu s}$
$1825\,\mathrm{\mu s}$
yes
Emission vs. $p/p_{\mathrm{sat}}$
Fig. 4
125 mT
$147\degree$
$500\,\mathrm{\mu s}$
$100\,\mathrm{\mu s}$
$70\,\mathrm{\mu s}$
$1825\,\mathrm{\mu s}$
yes
Dephasing
Fig. 5
125 mT
$147\degree$
$500\,\mathrm{\mu s}$
$100\,\mathrm{\mu s}$
$70\,\mathrm{\mu s}$
$1552\,\mathrm{\mu s}$
no
Scattering pattern
Fig. 9a
180 mT
$125\degree$
$10\,\mathrm{\mu s}$
$10\,\mathrm{\mu s}$
$122\,\mathrm{\mu s}$
no
$F_{r}$ vs. $p/p_{\mathrm{sat}}$
Fig. 10
125 mT
$147\degree$
$500\,\mathrm{\mu s}$
$100\,\mathrm{\mu s}$
$70\,\mathrm{\mu s}$
$1825\,\mathrm{\mu s}$
yes
CRC statistics
Fig. 12
125 mT
$147\degree$
$1000\,\mathrm{\mu s}$
$100\,\mathrm{\mu s}$
$50\,\mathrm{\mu s}$
$1825\,\mathrm{\mu s}$
yes
$F_{r}$ vs. $N_{c}$
Fig. 13
125 mT
$147\degree$
$1000\,\mathrm{\mu s}$
$100\,\mathrm{\mu s}$
$50\,\mathrm{\mu s}$
$1825\,\mathrm{\mu s}$
yes
CPMG-2
Fig. 14
125 mT
$147\degree$
$500\,\mathrm{\mu s}$
$100\,\mathrm{\mu s}$
$80\,\mathrm{\mu s}$
$3174\,\mathrm{\mu s}$
no
Quantum jumps
Fig. 15
125 mT
$147\degree$
$500\,\mathrm{\mu s}$
$200\,\mathrm{\mu s}$
$1000\,\mathrm{\mu s}$
$2123\,\mathrm{\mu s}$
yes

### v2

@@ C.3 Extended data
Figure 12: Timing diagrams of the experiments in: (a) Fig. 3, (b) Fig. 4b, and (c) Fig. 5. Note that (b) is a simplification because the waveform used is identical in timing to (a) including two readout steps and two CRCs; only the first of each is considered for Fig. 4b. Illustrations are not to scale. Steps are generally separated by $2\,\mathrm{\mu s}$ pauses (not shown). Weak measurement pulses in (b) and (c) resonantly drive the A1 transition, as in the spin polarization and projective measurement steps.
C.3.1 Charge resonance checks
@@
The readout characterization experiment in Fig. 3 includes charge resonance checks (CRCs) applied at both the beginning and end of each characterization cycle. A CRC consists of driving both the A1 and B2 transitions while simultaneously collecting emission. Many photons are collected during the CRC if the excitation is indeed resonant (e.g., neither the laser nor transition frequencies have drifted) and the system has not left the qubit subspace (e.g., it has not “blinked off”).
@@
The distribution of counts collected during a CRC are shown in Fig. 13, using data from the “wfm-A” experiment in Fig. 3. Data from wfm-B are nearly identical. The CRC result is recorded as a “pass” if $\geq N_{c}$ counts are measured and a “fail” if $<N_{c}$ counts are measured, where $N_{c}$ is the CRC discrimination threshold. Choosing $N_{c}=30$ for both CRCs for the data in Fig. 3, we find that $7.4\%$ of cycles of wfm-A pass both CRCs, and $7.7\%$ of cycles for wfm-B pass both CRCs, out of $4.27\times 10^{5}$ cycles total for each waveform.
@@
The probability that consecutive cycles pass either the first or second check is plotted in Fig. 13c, using a pass threshold of $N_{c}=6$ (the inflection threshold between the dark and bright distributions, Fig. 13b) and using a cycle time of $\tau_{\mathrm{cycle}}=1.825\,\mathrm{ms}$. Success probability of a given number of consecutive passes $n$ decays as $\propto e^{-n/N_{\mathrm{pass}}}$, consistent with fail events occurring at random. However, probability falls faster for the second check, which may result from drive-induced blinking [39].
[…]
@@
Figure 13: (a) Example of counts recorded during 400 CRCs. Each CRC consists of simultaneously driving the A1/B2 transitions. A threshold, here specified by the pink line at $N_{c}=30$, is used to discriminate between passing and failing checks. Data in the top (bottom) are from the first (second) check of the sequence used in Fig. 3 (see Fig. 12a for details), which is repeated $4.27\times 10^{5}$ times. (b) Distribution of counts collected during all CRCs in the Fig. 3 dataset. Separate distributions with few or many counts per check are separated by an inflection point around six counts. Choosing a pass threshold of $N_{c}=30$ leads to approximately $15\%$ cycles passing the first check, and, of these, approximately $50\%$ pass the second check. (c) Probability of consecutive passes before a fail is recorded, also using the data from Fig. 3 and using $N_{c}=6$. Data are fit to exponential decay. Faster decay is observed for the second check, perhaps due to drive-induced charging in the intervening readout steps. In all panels, only data from wfm-A is plotted; data for wfm-B is nearly identical.
@@
Figure 14: Analysis of the time-tagged dataset shown in Fig. 3, as a function of CRC threshold photon number $N_{c}$. If both CRCs per cycle each produce a photon number $\geq N_{c}$, the readout steps in that cycle are kept for analysis. (a) Readout fidelity $F_{r}$ and conditional fidelity $F_{c}$ (first readout step and wfm-A, respectively; results from the second readout step and wfm-B look nearly identical). (b) Mean photon number in the bright and dark distributions (first readout step; results from the second step look nearly identical). (c) Probability of passing each CRC, both CRCs, and the probability of the second CRC given a pass of the first. (Data are from wfm-A; results from wfm-B look nearly identical.)
C.3.2 Readout fidelity vs power
Figure 15: (a) Readout fidelity vs power. Data are fit to Eq. 28. (b) Fraction of CRCs passed, vs variable power of a readout step placed before the CRC. Both panels use the same dataset as Fig. 4.
Studying the power dependence of readout is important for the optimization of fidelity. Notice in Fig. 4b that, while the number of signal photons $\bar{n}_{b}-\bar{n}_{d}$ plateaus near saturation, the total number of collected photons continues to increase with power. We attribute this rise to scatter of the excitation laser into the collection path (see Appendix A.1 for details).
This power dependence makes clear that readout fidelity can be optimized by operating near saturation. Below saturation, increasing power simply produces more signal. Above saturation, however, driving with more power does not increase signal and can only cause rising noise. To test this intuition, we plot readout fidelity vs power in Fig. 15a, measured from the same dataset used to produce Fig. 4b. Data are fit to the model in Eq. 28. This model assumes that emitted photons follow Poissonian statistics; see Appendix A.1 for a derivation.
Note that this fit uses $p_{\mathrm{sat}}=313\,\mathrm{nW}$ and $\eta=0.992\times 10^{-3}$ obtained from fits to Fig. 4b, and fits for $f_{0}=0.89\pm 0.032$ and $\Lambda=977\pm 72$ as free parameters. Fixing $\Lambda=2244$ as a fixed parameter, as expected from the data in Fig. 2, does not produce a fit which closely matches the data in Fig. 4b. This discrepancy could be attributed to non-Poissonian statistics of light emitted from the SnV- (an assumption of Eq. 28), and/or from detrimental effects due to SnV- instability (e.g. cycles where the CRC is passed, but a blinking event occurs before or during the readout step). As such, the models in Eqs. 6 and 28 should be taken as simplifications that do not capture every possible effect on readout of the system, and the parameters reported from these fits should be taken as estimates based on these simplified assumptions only.
Finally, to verify the effect shown in Fig. 5 is truly measurement-induced dephasing, not, for example, drive-induced charging and/or heating, we study the fraction of CRCs passed as a function of weak measurement power. This is done using the same dataset as in Fig. 4b, and using a CRC at the end of the cycle after the weak measurement step. Results are show in Fig. 15b: As weak measurement power is increased, the fraction of passed checks changes little at $0.16\pm 0.03$, for all data except the highest measured power (at 19 times saturation) in which the fraction of passed checks drops to $0.09$. This suggests that the power-dependent increase in $\Gamma_{\phi}$ observed in Fig. 5 is not due to an effect which causes power-dependent blinking and/or spectral diffusion. This finding is consistent with prior demonstration of coherent spin control using an optical Raman drive [29], a technique which cannot work if a detuned laser causes significant error.
@@ C.3.3 Coherence time
The qubit’s coherence time is measured to be $T_{2}^{\mathrm{CPMG-2}}=270\pm 30\,\mathrm{\mu s}$ using a CPMG-2 pulse sequence [77] (Fig. 16) operating at $|\vec{B}|=125\,\mathrm{mT}$ and $\zeta=147\degree$. This measurement is comparable to results in Ref. [30], which operate at $\Lambda\approx 80$. For spin-1/2 group-IV qubits, at some magnetic field orientations, dynamical decoupling has a limited effect because the qubit’s $g$ factor is too similar to the predominantly spin-1/2 bath [30]. The measurement in Fig. 16 shows that here we avoid this detrimental effect.
@@ C.3.4 Quantum jumps
(c) The same time trace of the quantum jumps with assigned spin state as 0 or 1 based on the discrimination threshold ($N_{r}$).
We take advantage of single-shot readout to observe real-time quantum jumps of the electron spin [46]. The protocol, illustrated in Fig. 17a, consists of a green repump, then spin polarization, followed by readout on the same state that was polarized in the presence of a weak microwave drive. During this readout step, time-tagged photon detection events are binned into $5\,\mathrm{\mu s}$ windows and plotted as a time trace in Fig. 17b. In the absence of the weak spin drive, we expect zero photons due to high-fidelity initialization of the spin state and, thus, a complete dark time trace of the collected fluorescence. However, the weak spin drive (microwave power approximately $1\leavevmode\nobreak\ \,\mathrm{dBm}$ at the sample) occasionally kicks the spin state from $\ket{\uparrow}$ to $\ket{\downarrow}$, for example, when initialized using the A1 transition. This results in photons being emitted which can be seen as discrete spikes in the time trace as our readout determines the state in $\ket{\downarrow}$. Using a photon threshold of $N_{r}\geq 1$ (black dashed line), we distinguish between the bright and dark states and denote the electron being in the $\ket{\downarrow}$ ($\ket{\uparrow}$) state. Optical reinitialization also happens naturally under the continuous pump, so the jumping between the spin states (bright and dark) is a mixture of bidirectional mixing (microwave) and one-way reinitialization (bright to dark). To ensure capturing the quantum jumps is not convoluted with charge or spectral hopping, the entire sequence is bracketed with CRCs before and after each measurement shot (using a photon threshold of $N_{c}=30$ for both CRCs). In this experiment, the spin-flip process from $T_{1}$ (the order of hertz at the 1.7 K sample temperature [30, 31]) is much slower than the readout process. For clarity, Fig. 17c shows the same time trace where the electronic state is assigned as 0 or 1, based on the discrimination threshold ($N_{r}$).
C.3.5 Measurement-induced dephasing control experiment
Finally, we do a control experiment to demonstrate again that measurement-induced dephasing is indeed responsible for the dephasing shown in Fig. 5 rather than any alternative laser-induced process. This experiment is described in Fig. 18 and consists of comparing dephasing due to a resonant weak measurement pulse to dephasing caused by a control pulse that is off resonant only. No appreciable dephasing is observed during the control pulse, emphasizing that the dephasing observed in Fig. 5 is indeed caused by measurement.
Figure 18: Comparison between the measurement-induced dephasing experiment (as in Fig. 5, with pulse sequence in Fig. 12c) to a control. In (a) the experiment, the weak measurement pulse is applied by modulating both the laser amplitude with an acousto-optic modulator (AOM) and modulating an electro-optic modulator (EOM) whose first sideband (here at 3.718 GHz) is resonant with one of the SnV-’s spin-preserving optical transitions. (b) The control sequence is otherwise identical but does not modulate the EOM during the weak measurement pulse, so that all laser power remains in the detuned carrier. (c) Measurement-induced dephasing is observed only when resonantly driving the transition. Data are shown for a power of 1.3 nW entering the cryostat during the weak measurement pulse, causing a dephasing rate of $\Gamma_{\phi}/2\pi=104\pm 11\,\mathrm{kHz}$.
@@ C.4 List of data
Table 3: Summary of data and operating conditions. Repump power is typically the order of $100\,\mathrm{\mu W}$ and resonant excitation power is typically the order of $1-10\,\mathrm{\mu W}$, both specified going into the cryostat. Note that in all data the qubit is driven with the first sideband of the laser created by an electro-optical modulator, not the carrier signal.
Data Figure Magnet: amplitude $|\vec{B}|$ Magnet: angle $\zeta$ Repump duration Init. duration Readout duration Cycle time CRC?
Transitions vs $\zeta$ Fig. 2c 180 mT Swept $1\,\mathrm{\mu s}$ $1\,\mathrm{\mu s}$ $22\,\mathrm{\mu s}$ No
Spin polarization Fig. 2d 180 mT $53\degree$ and $147\degree$ $100\,\mathrm{\mu s}$ $100\,\mathrm{\mu s}$ $202\,\mathrm{\mu s}$ No
Cyclicity vs $\zeta$ Fig. 2e 180 mT Swept $100\,\mathrm{\mu s}$ $100\,\mathrm{\mu s}$ $202\,\mathrm{\mu s}$ No
Rabi Fig. 2f 125 mT $147\degree$ $500\,\mathrm{\mu s}$ $100\,\mathrm{\mu s}$ $100\,\mathrm{\mu s}$ $719\,\mathrm{\mu s}$ No
Rabi vs $\zeta$ Fig. 2g 132-183 mT $147\degree$ $20-500\,\mathrm{\mu s}$ $5-120\,\mathrm{\mu s}$ $1-60\,\mathrm{\mu s}$ $44-679\,\mathrm{\mu s}$ No
Single-shot readout Fig. 3 125 mT $147\degree$ $1000\,\mathrm{\mu s}$ $100\,\mathrm{\mu s}$ $50\,\mathrm{\mu s}$ $1825\,\mathrm{\mu s}$ Yes
Emission vs $p/p_{\mathrm{sat}}$ Fig. 4 125 mT $147\degree$ $500\,\mathrm{\mu s}$ $100\,\mathrm{\mu s}$ $70\,\mathrm{\mu s}$ $1825\,\mathrm{\mu s}$ Yes
Dephasing Fig. 5 125 mT $147\degree$ $500\,\mathrm{\mu s}$ $100\,\mathrm{\mu s}$ $70\,\mathrm{\mu s}$ $1552\,\mathrm{\mu s}$ No
Scattering pattern Fig. 11a 180 mT $125\degree$ $10\,\mathrm{\mu s}$ $10\,\mathrm{\mu s}$ $122\,\mathrm{\mu s}$ No
$F_{r}$ vs $p/p_{\mathrm{sat}}$ Fig. 15 125 mT $147\degree$ $500\,\mathrm{\mu s}$ $100\,\mathrm{\mu s}$ $70\,\mathrm{\mu s}$ $1825\,\mathrm{\mu s}$ Yes
CRC statistics Fig. 13 125 mT $147\degree$ $1000\,\mathrm{\mu s}$ $100\,\mathrm{\mu s}$ $50\,\mathrm{\mu s}$ $1825\,\mathrm{\mu s}$ Yes
$F_{r}$ vs $N_{c}$ Fig. 14 125 mT $147\degree$ $1000\,\mathrm{\mu s}$ $100\,\mathrm{\mu s}$ $50\,\mathrm{\mu s}$ $1825\,\mathrm{\mu s}$ Yes
CPMG-2 Fig. 16 125 mT $147\degree$ $500\,\mathrm{\mu s}$ $100\,\mathrm{\mu s}$ $80\,\mathrm{\mu s}$ $3174\,\mathrm{\mu s}$ No
Quantum jumps Fig. 17 125 mT $147\degree$ $500\,\mathrm{\mu s}$ $200\,\mathrm{\mu s}$ $1000\,\mathrm{\mu s}$ $2123\,\mathrm{\mu s}$ Yes
Dephasing control Fig. 18 125 mT $147\degree$ $500\,\mathrm{\mu s}$ $100\,\mathrm{\mu s}$ $50\,\mathrm{\mu s}$ $3433\,\mathrm{\mu s}$ No